OpenStudy (anonymous):

Assume f(x) = 2x + 8 and g(x) = 5. What is the value of (f + g)(3)?

OpenStudy (anonymous):

19?

OpenStudy (saifoo.khan):

(f+g) = 2x+8 + 5 (f+g) = 2x+8 + 5 (f+g)(3) = 2(3) + 13 = 6 + 13

OpenStudy (anonymous):

yep